# Online Detecting LLM-Generated Texts via Sequential Hypothesis Testing by Betting

### v2

@@ Abstract
Developing algorithms to differentiate between machine-generated texts and human-written texts has garnered substantial attention in recent years. Existing methods in this direction typically concern an offline setting where a dataset containing a mix of real and machine-generated texts is given upfront, and the task is to determine whether each sample in the dataset is from a large language model (LLM) or a human. However, in many practical scenarios, sources such as news websites, social media accounts, or on other forums publish content in a streaming fashion. Therefore, in this online scenario, how to quickly and accurately determine whether the source is an LLM with strong statistical guarantees is crucial for these media or platforms to function effectively and prevent the spread of misinformation and other potential misuse of LLMs. To tackle the problem of online detection, we develop an algorithm based on the techniques of sequential hypothesis testing by betting that not only builds upon and complements existing offline detection techniques but also enjoys statistical guarantees, which include a controlled false positive rate and the expected time to correctly identify a source as an LLM. Experiments were conducted to demonstrate the effectiveness of our method.
11footnotetext: The code and data are released at https://github.com/canchen-cc/online-llm-detection.
@@ 1 Introduction
While these existing methods can efficiently identify a text source in an offline setting where all the texts to be classified are provided in a single shot, they are not specifically designed to handle scenarios where texts arrive sequentially, and the goal is to quickly detect if the source is an LLM with strong statistical guarantees. Moreover, the process of most of the related works for offline detecting LLMs involves computing a score for each sample to determine how likely it is machine-generated and classify a text as a machine-generated one by determining whether the score passes a certain threshold. The threshold is a critical parameter that needs to be tuned using a validation dataset and must be determined before classifying testing data. Therefore, these offline detection methods fall short in the online scenarios, where one would like to design efficient algorithms with robust performance metrics, such as being a valid level-$\alpha$ test, having asymptotic power $1$, and enjoying a bound on the expected detection time for LLMs, which can be appealing for time-sensitive applications like monitoring online platforms.
@@
For example, the American Federal Communications Commission in 2017 decided to repeal net neutrality rules according to the public opinions collected through an online platform (Selyukh, 2017; Weiss, 2019). However, it was ultimately discovered that the overwhelming majority of the total $22$ million comments that support rescinding the rules were machine-generated (Kao, 2017). In 2019, Weiss (2019) used GPT-2 to overwhelm a website for collecting public comments on a medical reform waiver within only four days, where machine-generated comments eventually made up $55.3\%$ of all the comments (more precisely, $1,001$ out of $1,810$ comments). As discussed by Fröhling and Zubiaga (2021), a GPT-J model trained on a politics message board was then deployed on the same forum. It generated posts that included objectionable content and accounted for about $10\%$ of all activity during peak times (Kilcher, 2022). Furthermore, other online attacks mentioned by Fröhling and Zubiaga (2021) may even manipulate public discourse (Ferrara et al., 2016), flood news with fake content (Belz, 2019), or fraud by impersonating others on the Internet or via e-mail (Solaiman et al., 2019). However, existing bot detection methods for social media (e.g., Davis et al. (2016); Varol et al. (2017); Pozzana and Ferrara (2020); Ferrara (2023) and the references therein) might not be directly applicable to the online setting with strong statistical guarantees, to the best of our knowledge, and they often require training on extensive labeled datasets beforehand. This highlights the urgent need to develop algorithms with strong statistical guarantees that can quickly and effectively identify machine-generated texts in a timely manner, which has been largely overlooked in the literature.
@@
Our goal, therefore, is to tackle the problem of online detecting LLM-generated texts. More precisely, building upon existing score functions from those “offline approaches”, we aim to quickly determine whether the source of a sequence of texts observed in a streaming fashion is an LLM or a human, with a concrete bound on the required samples to detect LLMs and without the need for tuning a threshold of the affinity score that is typically used to classify data in those offline approaches. Our algorithm leverages the techniques of sequential hypothesis testing (Shafer, 2021; Ramdas et al., 2023; Shekhar and Ramdas, 2023). Specifically, we frame the problem of online detecting LLMs as a sequential hypothesis testing problem, where at each round $t$, a text from an unknown source is observed, and we aim to infer whether it is generated by an LLM. We also assume that a pool of examples of human-written texts is available, and our algorithm can sample a text from this pool of examples at any time $t$. Our method constructs a null hypothesis $H_{0}$ (to be elaborated soon), for which correctly rejecting the null hypothesis implies that the algorithm correctly identifies the source as an LLM under a mild assumption. Furthermore, since it is desirable to quickly identify an LLM when it is present and avoid erroneously declaring the source as an LLM, we also aim to control the type-I error (false positive rate) while maximizing the power to reduce type-II error (false negative rate), and to establish an upper bound on the expected total number of rounds to declare that the source is an LLM. We emphasize that our approach is non-parametric, and hence it does not need to assume that the underlying data of human or machine-generated texts follow a certain distribution (Balsubramani and Ramdas, 2015). It also avoids the need for assuming that the sample size is fixed or to specify it before the testing starts, and hence it is in contrast with some typical hypothesis testing methods that do not enjoy strong statistical guarantees in the anytime-valid fashion (Garson, 2012; Good, 2013; Tartakovsky et al., 2014). The way to achieve these is based on recent developments in sequential hypothesis testing via betting (Shafer, 2021; Shekhar and Ramdas, 2023). The setting of online testing with anytime-valid guarantees could be particularly useful when one seeks substantial savings in both data collection and time without compromising the reliability of their statistical testing. These desiderata might be elusive for approaches that based on collecting data in batch and classifying them offline to achieve. The effectiveness of our method is evaluated through comprehensive experiments.
@@
Figure 1: Overview of LLM Detection via Online Optimization. We sequentially observe text $y_{t}$ generated by an unknown source starting from time $t=1$ and aim to determine whether these texts are produced by a human or an LLM. The detection process can be divided into three steps. (1) (Score) At each time $t$, text $x_{t}$ and $y_{t}$ are evaluated by a selected score function $\phi(\cdot)$, where the sample $x_{t}$ is drawn from a prepared dataset consisting of human-written text examples. (2) (Update) The parameter $\theta_{t}$ is updated via the Online Newton Step (ONS) to increase the wealth $W_{t}$ rapidly when $y_{t}$ is an LLM-generated text. A large value of $W_{t}$ serves as significant evidence and provides confidence to declare that the unknown source is an LLM. (3) (Check) Whether the wealth $W_{t}\geq 1/\alpha$ is checked. If this event happens, we declare the unknown source of $y_{t}$ as LLM. Otherwise, if the time budget $T$ is not yet exhausted or if we have an unlimited time budget, we proceed to $t+1$ and repeat the steps. When $t=T$, if the condition $W_{T}\geq Z/\alpha$ holds, where $Z$ is drawn from an uniform distribution in $[0,1]$, our algorithm will also declare the source as an LLM.
@@ 2 Preliminaries
Sequential Hypothesis Test with Level-$\alpha$ and Asymptotic Power One. Let us denote a forward filtration $\mathcal{F}=(\mathcal{F}_{t})_{t\geq 0}$, where $\mathcal{F}_{t}=\sigma(Z_{1},\ldots,Z_{t})$ represents an increasing sequence that accumulates all the information from the observations $\{Z_{i}:i\geq 1\}$ up to time point $t$. A process $W:=(W_{t})_{t\geq 1}$, adapted to $(\mathcal{F}_{t})_{t\geq 1}$, is defined as a P-martingale if it satisfies $E_{P}[W_{t}|\mathcal{F}_{t-1}]=W_{t-1}$ for all $t\geq 1$. Furthermore, $W$ is a P-supermartingale if $E_{P}[W_{t}|\mathcal{F}_{t-1}]\leq W_{t-1}$ for all $t\geq 1$. In our algorithm design, we will consider a martingale $W$ and define the event $\{W_{t}\geq{1}/{\alpha}\}$ as rejecting the null hypothesis $H_{0}$, where $\alpha>0$ is a user-specified “significance level” parameter. We denote the stopping time $\tau:=\arg\inf_{t\geq 1}P\{W_{t}\geq{1}/{\alpha}\}$ accordingly.
We further recall that a hypothesis test is a level-$\alpha$ test if $\sup_{P\in H_{0}}P\left(\exists t\geq 1,W_{t}\geq{1}/{\alpha}\right)\leq\alpha$, or alternatively, if $\sup_{P\in H_{0}}P\left(\tau\leq\infty\right)\leq\alpha.$ Furthermore, a test has asymptotic power $(1-\beta)$ if $\sup_{P\in H_{1}}P\left(\forall t\geq 1,W_{t}<{1}/{\alpha}\right)\leq\beta$, or if $\sup_{P\in H_{1}}P\left(\tau=\infty\right)\leq\beta$, where $H_{1}$ represents the alternative hypothesis. A test with asymptotic power one (i.e., $\beta=0$) means that when the alternative hypothesis is true, the test will eventually rejects the null hypothesis $H_{0}$. As shown later, our algorithm that will be introduced shortly is a provable sequential hypothesis testing method with level-$\alpha$ and asymptotic power one.
Problem Setup. We consider a scenario in which, at each round $t$, a text $y_{t}$ from an unknown source is observed, and additionally, a human-written text $x_{t}$ can be sampled from a dataset of human-written text examples at our disposal. The goal is to quickly and correctly determine whether the source that produces the sequence of texts $\{y_{t}\}_{t=1}^{T}$ is an LLM or a human. We assume that a score function $\phi(\cdot):\text{Text}\rightarrow\mathbb{R}$ is available, which, given a text as input, outputs a score. The score function $\phi(\cdot)$ that we consider in this work are those proposed for detecting LLM-generated texts in offline settings, e.g., Mitchell et al. (2023); Bao et al. (2023); Su et al. (2023); Bao et al. (2023); Yang et al. (2023). We provide more details on these score functions in the experiments section and in the exposition of the literature in Appendix A.1.
@@
We note that when $H_{0}$ is true, this is not equivalent to saying that the texts $\{y_{t}\}_{t=1}^{T}$ are human-written, as different distributions can share the same mean. However, under the additional assumption of the existence of a good score function $\phi(\cdot)$ which produces scores for machine-generated texts with a mean $\mu_{y}$ different from that of human-generated texts $\mu_{x}$, $H_{0}$ is equivalent to the unknown source being human. Therefore, under this additional assumption, when the unknown source $\rho^{y}$ is an LLM, then rejecting the null hypothesis $H_{0}$ is equivalent to correctly identifying that the source is indeed an LLM. In our experiments, we found that this assumption generally holds empirically for the score functions that we adopt. That is, the empirical mean of $\phi(y_{t})$ significantly differs from that of $\phi(x_{t})$ when each $y_{t}$ is generated by an LLM. Figure 1 illustrates the process of online detecting LLMs.
@@
To connect the learner’s game with sequential hypothesis testing, one of the key techniques that will be used in the algorithm design and analysis is Ville’s inequality (Ville, 1939), which states that if $(W_{t})_{t\geq 1}$ is a nonnegative supermartingale, then one has $P(\exists t:W_{t}\geq{1}/{\alpha})\leq\alpha\mathbb{E}[W_{0}]$. The idea is that if we can guarantee that the learner’s wealth $W_{t}$ can remain nonnegative from $W_{0}=1$, then Ville’s inequality can be used to control the type-I error at level $\alpha$ simultaneously at all time steps $t$. To see this, let $g_{t}=\phi({x}_{t})-\phi({y}_{t})$. Then, when $P\in H_{0}$ (i.e., the null hypothesis $\mu_{x}=\mu_{y}$ holds), the wealth $(W_{t})_{t\geq 1}$ is a P-supermartingale, because
$\mathbb{E}_{P}[W_{t}|F_{t-1}]=\mathbb{E}_{P}[W_{t-1}(1-\theta_{t}g_{t})|F_{t-1}]=W_{t-1}\cdot\mathbb{E}_{P}[(1-\theta_{t}(\phi({x}_{t})-\phi({y}_{t})))|F_{t-1}]=W_{t-1}.$
@@
Algorithm 1 Online Detecting LLMs via Online Optimization and Betting
@@
3:# $T$ is the time budget, which can be $\infty$ if their is no time constraint.
@@
13:     Specify the decision space $\mathcal{K}_{t+1}:=[-\frac{1}{2d_{t+1}},\frac{1}{2d_{t+1}}]$ to ensure $W_{t+1}$ is nonnegative.
14:     // Update $\theta_{t+1}\in\mathcal{K}_{t+1}$ via $\mathrm{ONS}$ on the loss function $\ell_{t}(\theta):=-\ln(1-g_{t}\theta)$.
15:     Compute $z_{t}=\frac{d\ell_{t}(\theta_{t})}{d\theta}=\frac{g_{t}}{1-g_{t}\theta_{t}}$ and $a_{t}=a_{t-1}+z_{t}^{2}$.
16:     Compute $\theta_{t+1}=\max\left(\min\left(\theta_{t}-\frac{1}{\gamma}\frac{z_{t}}{a_{t}},\frac{1}{2d_{t+1}}\right),-\frac{1}{2d_{t+1}}\right)$.

### v3

# Online Detection of LLM-Generated Texts via Sequential Hypothesis Testing by Betting

@@ Abstract
Developing algorithms to differentiate between machine-generated texts and human-written texts has garnered substantial attention in recent years. Existing methods in this direction typically concern an offline setting where a dataset containing a mix of real and machine-generated texts is given upfront, and the task is to determine whether each sample in the dataset is from a large language model (LLM) or a human. However, in many practical scenarios, sources such as news websites, social media accounts, and online forums publish content in a streaming fashion. Therefore, in this online scenario, how to quickly and accurately determine whether the source is an LLM with strong statistical guarantees is crucial for these media or platforms to function effectively and prevent the spread of misinformation and other potential misuse of LLMs. To tackle the problem of online detection, we develop an algorithm based on the techniques of sequential hypothesis testing by betting that not only builds upon and complements existing offline detection techniques but also enjoys statistical guarantees, which include a controlled false positive rate and the expected time to correctly identify a source as an LLM. Experiments were conducted to demonstrate the effectiveness of our method.
@@ 1 Introduction
While these existing methods can efficiently identify a text source in an offline setting where each text is classified independently, they are not specifically designed to handle scenarios where texts arrive sequentially, and the goal is to quickly detect if the source is an LLM with strong statistical guarantees. A common limitation of most detectors is that their offline detection process involves computing a score for each sample to determine how likely it is machine-generated and classify a text as a machine-generated one by determining whether the score passes a certain threshold. The threshold is a critical parameter that needs to be tuned using a validation dataset and must be determined before classifying testing data. Moreover, existing offline detectors are unable to control the type-I error rate (false positive rate) in an online setting—even for state-of-the-art zero-shot detectors such as Fast-DetectGPT (Bao et al., 2023), Binoculars (Hans et al., 2024), or other efficient models like RoBERTa-Base/Large (Liu et al., 2019), ReMoDetect (Lee et al., 2024), and Raidar (Mao et al., 2024), which do not require threshold tuning at inference time.
For example, a naive way to adapt an offline detector to an online scenario is to declare the source as an LLM as soon as any single text is predicted as such. However, this adaptation leads to a false positive rate that converges to $1$ as the number of texts becomes sufficiently large—unless the detector is perfectly ($100\%$) accurate on human-written texts. Formally, if the detector has accuracy $\delta$ on human-written texts, then under the naive reduction approach to online detection, the false positive rate over $T$ rounds will be $1-\delta^{T}$, which approaches $1$ as $T$ grows, even when $\delta$ is very close to $1$.
@@
Therefore, these offline detection methods fall short in the online scenarios, where one would like to design efficient algorithms with robust performance metrics, such as being a valid level-$\alpha$ test, having asymptotic power $1$, and enjoying a bound on the expected detection time for LLMs, which can be appealing for time-sensitive applications like monitoring online platforms. For example, the American Federal Communications Commission in 2017 decided to repeal net neutrality rules according to the public opinions collected through an online platform (Selyukh, 2017; Weiss, 2019). However, it was ultimately discovered that the overwhelming majority of the total $22$ million comments that support rescinding the rules were machine-generated (Kao, 2017). In 2019, Weiss (2019) used GPT-2 to overwhelm a website for collecting public comments on a medical reform waiver within only four days, where machine-generated comments eventually made up $55.3\%$ of all the comments (more precisely, $1,001$ out of $1,810$ comments). As discussed by Fröhling and Zubiaga (2021), a GPT-J model trained on a politics message board was then deployed on the same forum. It generated posts that included objectionable content and accounted for about $10\%$ of all activity during peak times (Kilcher, 2022). Furthermore, other online attacks mentioned by Fröhling and Zubiaga (2021) may even manipulate public discourse (Ferrara et al., 2016), flood news with fake content (Belz, 2019), or fraud by impersonating others on the Internet or via e-mail (Solaiman et al., 2019). However, to the best of our knowledge, existing bot detection methods for social media (e.g., Davis et al. (2016); Varol et al. (2017); Pozzana and Ferrara (2020); Ferrara (2023) and the references therein) might not be directly applicable to the online setting with strong statistical guarantees, and they often require training on extensive labeled datasets beforehand. This highlights the urgent need to develop algorithms with strong statistical guarantees that can quickly and effectively identify machine-generated texts in a timely manner, which has been largely overlooked in the literature.
@@
Our goal, therefore, is to tackle the problem of online detection of LLM-generated texts. More precisely, building upon existing score functions from those “offline approaches”, we aim to quickly determine whether the source of a sequence of texts observed in a streaming fashion is an LLM or a human, with a concrete bound on the required samples to detect LLMs and without the need for tuning a threshold of the affinity score that is typically used to classify data in those offline approaches. Our algorithm leverages the techniques of sequential hypothesis testing (Shafer, 2021; Ramdas et al., 2023; Shekhar and Ramdas, 2023). Specifically, we frame the problem of online LLM detection as a sequential hypothesis testing problem, where at each round $t$, a text from an unknown source is observed, and we aim to infer whether it is generated by an LLM. We also assume that a pool of examples of human-written texts is available, and our algorithm can sample a text from this pool of examples at any time $t$. Our method constructs a null hypothesis $H_{0}$ (to be elaborated soon), for which correctly rejecting the null hypothesis implies that the algorithm correctly identifies the source as an LLM under a mild assumption. Furthermore, since it is desirable to quickly identify an LLM when it is present and avoid erroneously declaring the source as an LLM, we also aim to control the type-I error rate (false positive rate) while maximizing the power to reduce type-II error rate (false negative rate), and to establish an upper bound on the expected total number of rounds to declare that the source is an LLM. We emphasize that our approach is non-parametric, and hence it does not need to assume that the underlying data of human or machine-generated texts follow a certain distribution (Balsubramani and Ramdas, 2015). It also avoids the need for assuming that the sample size is fixed or to specify it before the testing starts, and hence it is in contrast with some typical hypothesis testing methods that do not enjoy strong statistical guarantees in the anytime-valid fashion (Garson, 2012; Good, 2013; Tartakovsky et al., 2014). The way to achieve these is based on recent developments in sequential hypothesis testing via betting (Shafer, 2021; Shekhar and Ramdas, 2023). The setting of online testing with anytime-valid guarantees could be particularly useful when one seeks substantial savings in both data collection and time without compromising the reliability of their statistical testing. These desiderata might be elusive for approaches that based on collecting data in batch and classifying them offline to achieve.
We evaluate the effectiveness of our method through comprehensive experiments. The code and datasets are available via this link: https://github.com/canchen-cc/online-llm-detection.
@@
Figure 1: Overview of LLM Detection via Online Optimization. (0) (Observe) We sequentially observe text $y_{t}$ generated by an unknown source starting from time $t=1$ and aim to determine whether these texts are produced by a human or an LLM. The detection process can be divided into three steps. (1) (Score) At each time $t$, text $x_{t}$ and $y_{t}$ are evaluated by a selected score function $\phi(\cdot)$, where the sample $x_{t}$ is drawn from a prepared dataset consisting of human-written text examples. (2) (Update) The parameter $\theta_{t}$ is updated via the Online Newton Step (ONS) to increase the wealth $W_{t}$ rapidly when $y_{t}$ is an LLM-generated text. A large value of $W_{t}$ serves as significant evidence and provides confidence to declare that the unknown source is an LLM. (3) (Check) Whether the wealth $W_{t}\geq 1/\alpha$ is checked. If this event happens, we declare the unknown source of $y_{t}$ as LLM. Otherwise, if the time budget $T$ is not yet exhausted or if we have an unlimited time budget, we proceed to $t+1$ and repeat the steps. When $t=T$, if the condition $W_{T}\geq Z/\alpha$ holds, where $Z$ is drawn from an uniform distribution in $[0,1]$, our algorithm will also declare the source as an LLM.
@@ 2 Preliminaries
Sequential Hypothesis Testing with Level-$\alpha$ and Asymptotic Power One. Let us denote a forward filtration $\mathcal{F}=(\mathcal{F}_{t})_{t\geq 0}$, where $\mathcal{F}_{t}=\sigma(Z_{1},\ldots,Z_{t})$ represents an increasing sequence that accumulates all the information from the observations $\{Z_{i}:i\geq 1\}$ up to time point $t$. A process $W:=(W_{t})_{t\geq 1}$, adapted to $(\mathcal{F}_{t})_{t\geq 1}$, is defined as a P-martingale if it satisfies $E_{P}[W_{t}|\mathcal{F}_{t-1}]=W_{t-1}$ for all $t\geq 1$. Furthermore, $W$ is a P-supermartingale if $E_{P}[W_{t}|\mathcal{F}_{t-1}]\leq W_{t-1}$ for all $t\geq 1$. In our algorithm design, we will consider a martingale $W$ and define the event $\{W_{t}\geq{1}/{\alpha}\}$ as rejecting the null hypothesis $H_{0}$, where $\alpha>0$ is a user-specified “significance level” parameter. We denote the stopping time $\tau:=\inf\{t\geq 1:W_{t}\geq\tfrac{1}{\alpha}\}$ accordingly.
We further recall that a hypothesis test is a level-$\alpha$ test if $\sup_{P\in H_{0}}P\left(\exists t\geq 1,W_{t}\geq{1}/{\alpha}\right)\leq\alpha$, or alternatively, if $\sup_{P\in H_{0}}P\left(\tau<\infty\right)\leq\alpha.$ Furthermore, a test has asymptotic power $1-\beta$ if $\sup_{P\in H_{1}}P\left(\forall t\geq 1,W_{t}<{1}/{\alpha}\right)\leq\beta$, or if $\sup_{P\in H_{1}}P\left(\tau=\infty\right)\leq\beta$, where $H_{1}$ represents the alternative hypothesis. A test with asymptotic power one (i.e., $\beta=0$) means that when the alternative hypothesis $H_{1}$ is true, the test will eventually rejects the null hypothesis $H_{0}$. As shown later, our algorithm that will be introduced shortly is a provable sequential hypothesis testing method with level-$\alpha$ and asymptotic power one.
Problem Setup. We consider a scenario in which, at each round $t$, a text $y_{t}$ from an unknown source is observed, and additionally, a human-written text $x_{t}$ can be sampled from a dataset of human-written text samples at our disposal. The goal is to quickly and correctly determine whether the source that produces the sequence of texts $\{y_{t}\}_{t=1}^{T}$ is an LLM or a human. We assume that a score function $\phi(\cdot):\text{Text}\rightarrow\mathbb{R}$ is available, which, given a text as input, outputs a score. The score function $\phi(\cdot)$ that we consider in this work are those proposed for detecting LLM-generated texts in offline settings, e.g., Mitchell et al. (2023); Bao et al. (2023); Su et al. (2023); Bao et al. (2023); Yang et al. (2023). We provide more details on these score functions in the experiments section and in the exposition of the literature in Appendix A.1.
@@
We note that when $H_{0}$ is true, this is not equivalent to saying that the texts $\{y_{t}\}_{t=1}^{T}$ are human-written, as different distributions can share the same mean. However, under the additional assumption of the existence of a good score function $\phi(\cdot)$ which produces scores for machine-generated texts with a mean $\mu_{y}$ different from that of human-generated texts $\mu_{x}$, $H_{0}$ is equivalent to the unknown source being human. Therefore, under this additional assumption, when the unknown source $\rho^{y}$ is an LLM, then rejecting the null hypothesis $H_{0}$ is equivalent to correctly identifying that the source is indeed an LLM. In our experiments, we found that this assumption holds empirically for the score functions we adopt. That is, the empirical mean of $\phi(y_{t})$ significantly differs from that of $\phi(x_{t})$ when each $y_{t}$ is generated by an LLM. Figure 1 illustrates the online detection process for LLM-generated texts.
@@
To connect the learner’s game with sequential hypothesis testing, one of the key techniques that will be used in the algorithm design and analysis is Ville’s inequality (Ville, 1939), which states that if $(W_{t})_{t\geq 1}$ is a nonnegative supermartingale, then one has $P(\exists t:W_{t}\geq{1}/{\alpha})\leq\alpha\mathbb{E}[W_{0}]$. The idea is that if we can guarantee that the learner’s wealth $W_{t}$ remains nonnegative from $W_{0}=1$, then Ville’s inequality can be used to control the type-I error at level $\alpha$ simultaneously at all time steps $t$. To see this, let $g_{t}=\phi({x}_{t})-\phi({y}_{t})$. Then, when $P\in H_{0}$ (i.e., the null hypothesis $\mu_{x}=\mu_{y}$ holds), the wealth $(W_{t})_{t\geq 1}$ is a P-supermartingale, because
$\mathbb{E}_{P}[W_{t}|F_{t-1}]=\mathbb{E}_{P}[W_{t-1}(1-\theta_{t}g_{t})|F_{t-1}]$
$=\mathbb{E}_{P}[W_{t-1}\cdot(1-\theta_{t}(\phi({x}_{t})-\phi({y}_{t})))|F_{t-1}]=W_{t-1}.$
@@
Algorithm 1 Online Detection of LLMs via Online Optimization and Betting
@@
3:# $T$ is the time budget, which can be set to $\infty$ if there is no time constraint.
@@
13:      Specify the decision space $\mathcal{K}_{t+1}$ as $\mathcal{K}_{t+1}:=[-\frac{1}{2d_{t+1}},\frac{1}{2d_{t+1}}]$ to ensure $W_{t+1}\geq 0$.
14:      # Update $\theta_{t+1}\in\mathcal{K}_{t+1}$ via $\mathrm{ONS}$ on the loss function $\ell_{t}(\theta):=-\ln(1-g_{t}\theta)$.
15:     Compute $z_{t}=\frac{d\ell_{t}(\theta_{t})}{d\theta}=\frac{g_{t}}{1-g_{t}\theta_{t}}$, $a_{t}=a_{t-1}+z_{t}^{2}$.
16:     Update $\theta_{t+1}=\max\left(\min\left(\theta_{t}-\frac{1}{\gamma}\frac{z_{t}}{a_{t}},\frac{1}{2d_{t+1}}\right),-\frac{1}{2d_{t+1}}\right)$.